OpenStudy (anonymous):

Will give medal to who ever answers first. Solve this problem using Elimination and Substitution, 7X-3X=20 5X+3Y=16 Please type it out.

OpenStudy (anonymous):

Simplifying 7x = 3x + 20 Reorder the terms: 7x = 20 + 3x Solving 7x = 20 + 3x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-3x' to each side of the equation. 7x + -3x = 20 + 3x + -3x Combine like terms: 7x + -3x = 4x 4x = 20 + 3x + -3x Combine like terms: 3x + -3x = 0 4x = 20 + 0 4x = 20 Divide each side by '4'. x = 5 Simplifying x = 5

OpenStudy (anonymous):

Simplifying 5x + 3y = 16 Solving 5x + 3y = 16 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-3y' to each side of the equation. 5x + 3y + -3y = 16 + -3y Combine like terms: 3y + -3y = 0 5x + 0 = 16 + -3y 5x = 16 + -3y Divide each side by '5'. x = 3.2 + -0.6y Simplifying x = 3.2 + -0.6y

OpenStudy (anonymous):

The X coordinates should be 3 and the y should be 1/3.

OpenStudy (anonymous):

what do you mean this isn't multiple choice is it

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I have to solve the problems for the x and y values.